OpenStudy (anonymous):

simplify forty two with the variable a subtracted by eighteen with a variable b, over six...

jigglypuff314 (jigglypuff314):

you mean like this? \[\frac{ 42a - 18b }{ 6 }\]

OpenStudy (anonymous):

it has a math processing error

jigglypuff314 (jigglypuff314):

42a-18b _________ 6 or 42a - (18b)/6

jigglypuff314 (jigglypuff314):

okay, so what do you get when you factor out a 6 from 42a - 18b =

jigglypuff314 (jigglypuff314):

42a - 18b = 6(7a) - 6(3b) = 6 (7a - 3b) 6/6 = 1 so you are left with 7a - 3b

OpenStudy (anonymous):

thank you so much! :)

jigglypuff314 (jigglypuff314):

you're welcome :)